OpenStudy (anonymous):

If you invested a penny on Jan 1, 1776 at 10% interest compounded daily, how much would you have on Jan 1, 2011 ?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(nt) A = 0.01(1+0.10/365)^(365*235) A = 160,148,364.838484 A = 160,148,364.84 So you would have $160,148,364.84 Note: n = 365 assuming there are 365 days in a year and t = 235 because 235 years elapsed from 1776 to 2011